. If a line is perpendicular to the line 5x - y = 0 and forms a triangle with coordinate axes of area 5 sq. units, then its equation is ____________________.

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
y=-1/5x+c

y/c+x/5c=1
.5*c*5c=5
c^2=2
c=sqrt(2)

the line is, y=.-2x+sqrt(2)

Sher Mohammad
B.Tech, IIT Delhi